# Quantum Entanglement without nonlocal causation in (3,2)-dimensional spacetime

### v2

@@ Abstract
This work aims at exploring whether the nonlocal correlations due to quantum entanglement could exist without nonlocal causation. This is done with the aid of a toy model to investigate whether the ability of two quantum entangled particles to ”correlate” their behaviors even at very large distances and in the absence of any physical connection can be seen as due to an exchange of information through an extra-temporal dimension. Since superluminal information exchange is obviously forbidden in our (3,1) space-time, an extra-temporal dimension is needed in order to recover the physical picture of finite velocity information exchange between entangled entities. Assuming that the geometry of space-time of dimension (3,2) is described by a metric containing a warping factor, the confinement of the massive particles in the extra time dimension follows. Therefore, why we do not experience an infinitely large extra time dimension can be explained. The toy model proposed here is defined by borrowing Bohm-Bub’s proposal to describe the wavefunction collapse by means of nonlinear (non-unitary) dynamical equations and then elaborating this approach for an entangled system. The model so obtained is just speculative without any claim of being robust against any criticism, nevertheless, it satisfies the purpose of giving the possibility to the hypotheses formulated above to be verified experimentally; in fact, it makes the proposal of an experiment potentially interesting which would otherwise be immediately dismissed as manifestly trivial. The proposed experiment would consist of checking the possible violation of Bell’s inequality between two identical but independent systems under appropriate conditions. Beyond its theoretical interest, entanglement is a key topic in quantum computing and quantum technologies, so any attempt to gain a deeper understanding of it could be useful.
@@ I Introduction
Then a strict and undeniable violation of Bell’s inequality was given by Alain Aspect’s experiments aspect1 ; aspect2 where each single channel polarizer was replaced by a fast switching device redirecting the incident light beam to a differently oriented polarizer. The detection events on opposite sides of the experimental apparatus were separated by a space-like interval, thus free of locality loopholes. Further confirmations of the violation of Bell’s inequality were then obtained at increasingly large spatial separations nl1 ; nl2 ; nl3 ; nl4 . Therefore, quantum nonlocality is undeniably a property of our physical world, but Bohr’s viewpoint can sound tautologic and unsatisfactory compared to our description of physical phenomena based on events located in space and evolving in time, so that the existence of correlations between events separated by an arbitrarily large space-like interval is a conundrum that goes against our intuitive perception of the physical world. Of course, not only quantum mechanics compels us to give up our intuitive perception of the physical world, nevertheless we can wonder whether - to some extent - we can recover a ”causal” description of quantum nonlocality by paying a different price with respect to Bohr’s claim of indivisibility of quantum systems.
The price to pay in the framework proposed in the present work is to consider the existence of an extra time dimension. In what follows there is no pretence of proposing something that could qualify as a new theory, rather, by means of a toy model we want to move some first steps in exploring whether nonlocality can be given a ”more familiar” interpretation by considering that while performing the measurement of an entangled state this appears as actually composed of independently existing units that exchange information via a sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ that propagates at finite velocity through an extra dimension of temporal kind $\tau$. This idea can be tested against an experiment that, if proved feasible and produced the hypothesized result, would make worthwhile and necessary to go beyond the toy model discussed below.
@@ I.1 A quick digression on extra time-dimensions
On the other hand, any formulation of fundamental physics with multiple times is somehow non-trivial. In fact, as clearly summarized in Ref.B3 , naive attempts to add extra time dimensions to existing frameworks have led to two major kinds of discouraging and seemingly unavoidable difficulties: the appearance of ghosts and violations of causality. Ghosts are quantum states of systems occurring with negative probability. Causality violations are exemplified by the so called ”grandfather paradox”: two-dimensions of time by allowing time travels would make possible to kill one’s ancestors before having being born, thus entailing an absurdity. This notwithstanding, an extensive work (an excerpt of which can be found in Refs.bars1 ; bars2 ; bars3 ; bars4 ; bars5 ; bars6 ) by Itzhak Bars and co-workers on a new gauge (symplectic) symmetry, called $Sp(2,{R})$, overcomes the mentioned problems and uniquely leads to the formalism of 2T-physics. In this two time theory the additional time dimension, being treated as a ”gauge” and implied by the non-triviality of the new gauge principle, seems to some extent unphysical.
@@
The paper is organized as follows. In Section II.1 we sketch the Bohm-Bub theory, then in Section II.2 we propose an extension of this theory to simultaneous measurements of a Bell state. In Section II.3 we suggest that the fundamental idea proposed in the present work could be tested against an experiment. Section III contains some concluding remarks. Finally, in the Appendix it is sketchily shown how the formulation of the above mentioned exotic KKTs can be borrowed to define a warped metric of the (3,2)-dimensional spacetime yielding the confinement of massive particles in the extra time dimension, thus explaining why we do not experience an infinitely large extra time dimension.
@@ II Nonlinear equations for wavefunction collapse
As is well known, two postulates of quantum mechanics are somewhat conflicting because on the one side the time evolution of the state vector of a given system is described by a linear, unitary operator and, on the other side, the result of the measure of an observable projects the system into the subspace relative to the eigenvalue/eigenstate of the measured observable according to the result obtained. Thus, measuring an observable entails the so-called collapse of the state vector (or wavefunction collapse), a discontinuous breach of the unitary evolution of a quantum system, actually a nonlinear time evolution. This topic has given rise to several discussions and interpretations (see for instance Ref.bassi ), however, for our purpose we resort to the proposal described in the following section.
@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter our main hypothesis enters the game. We assume that the wavefunctions collapses of the two particles - that now have their own individuality - are correlated via an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real physical field ${\mathscr{X}}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. We tentatively assume that the dynamical evolution of the scalar field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the field ${\mathscr{X}}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the outcome of the measurement unpredictable.
@@ II.3 Proposal for a nonconventional experiment
The source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$ in equation (20), even if of unspecified analytic form, amounts to assuming that the field ${\mathscr{X}}({\bf x},t,\tau)$ encodes the information about the wavefunction collapse due to the measurement performed on a given system, and equation (20) tells that this information is spread in every spatial direction. Therefore, a collapse information could also reach another identical system driving its collapse and mimicking entanglement correlation. This is the hypothesis to be experimentally proved or disproved that also suggests how to design an appropriate experiment. In fact, let us now consider two identical systems, depicted in Figure 3, where Alice and Bob perform measurements at the ends of the first system, and Eve and Tom at the ends of the second system. Of course, the photons (or particles with spin) reaching Alice and Bob or Eve and Tom are in state superposition and constrained to fulfil a conservation law (momentum, energy, angular momentum) which is not the case of two particles reaching, say, Alice and Eve because a-priori the entities (say photons, as in Figure 3) reaching Alice and Eve are independent, that is, not entangled. Now, as soon as the photon detected by Alice has ”chosen” its polarization state how does the photon being detected by Bob ”know” that it has no choice left for its polarization state? The orthodox reply is that they are part of the same indivisible system even if they are very far apart, but if we assume that the information about the outcome of the measurement performed by Alice is broadcasted in every spatial direction (through the extra temporal dimension), a particle belonging to an identical system, like the one reaching Eve, could be driven to collapse into the same complementary state as in Bob’s case. Therefore, let us imagine an experimental setup where the measurement apparatuses of Alice and Eve are spatially close to one another and both of them much farther from the apparatuses of their companions Bob and Tom, so that the extra-time interval $\Delta_{1}\tau$ needed by the hypothetical field ${\mathscr{X}}({\bf x},t,\tau)$ to exchange information between Alice and Bob on one side, and Eve and Tom on the other side, is much longer than the interval $\Delta_{2}\tau$ needed to exchange information between Alice and Eve. Consequently, the central hypothesis of the present work boils down to rewriting the collapse equations (II.2) as
@@
A clarification needs to be made here. Equations (II.2) and (II.3) describe the evolution of wavefunctions collapse in ordinary time after $t_{0}$, the instant of particles interaction with the measuring devices, and ”after” a mutual exchange of information in the extra time. In the case of the proposed new experiment, the surmised phenomenon of information exchange through the extra time depends entirely on the spatial arrangement of the four measuring devices. The exchange of information between Eve and Alice (assumed at the shortest distance among the four measuring devices) could conceivably destroy the entanglement with their respective partners allowing a rearrangement of the correlations among the results of the four measuring devices. A more refined analysis would be needed to describe what kind of correlations - among all these devices - could be expected with generic spatial arrangements. Such analysis is beyond the scope of our toy model and would be motivated by a positive experimental outcome of the proposed experiment outlined above.
@@ III Concluding remarks
It is not out of place to sketchily show how we can borrow from Ref.visser - almost verbatim - the suggestion for a five dimensional space-time metric of (+ + + - -) signature - with a non-compact and infinite extra time dimension - where the particles are trapped on the standard four-dimensional space-time. Thus, by simply modifying the signature of the extra dimension of the ”exotic” Kaluza-Klein metric of Ref.visser , consider the infinitesimal arc-length in 3+2 space-time as
@@
which has the form a ultrahyperbolic equation weinstein for which a solution can be found in the form
@@ Acknowledgments
The author wishes to thank Roger Penrose for an interesting and useful discussion held at the Arts Centre De Brakke Grond, Amsterdam, in 2014, a discussion that has been the remote origin of the present work. Useful comments and suggestions emerged during several discussions with Giulio Pettini, Gabriele Vezzosi, Matteo Gori, Roberto Franzosi, Guglielmo Iacomelli, and Jack Tuszynski. The author thanks Stefano Ruffo for having brought to his attention the paper in Ref.baracca .

### v3

@@ Abstract
This work aims at exploring whether the nonlocal correlations due to quantum entanglement could exist without nonlocal causation. This is done with the aid of a toy model to investigate whether the ability of two quantum entangled particles to ”correlate” their behaviors even at very large distances and in the absence of any physical connection can be seen as due to an exchange of information through an extra-temporal dimension. Since superluminal information exchange is forbidden in our (3,1) space-time, an extra-temporal dimension is needed to recover the physical picture of finite velocity information exchange between entangled entities. Assuming that the geometry of space-time of dimension (3,2) is described by a metric containing a warping factor, the confinement of the massive particles in the extra time dimension follows. Therefore, why we do not experience an infinitely large extra time dimension can be explained. The toy model proposed here is defined by borrowing Bohm-Bub’s proposal to describe the wavefunction collapse using nonlinear (non-unitary) dynamical equations and then elaborating this approach for an entangled system. The model so obtained is just speculative without any claim of being robust against any criticism, nevertheless, it satisfies the purpose of giving the possibility to the hypotheses formulated above to be verified experimentally; in fact, it proposes an experiment potentially interesting which would otherwise be immediately dismissed as manifestly trivial. The proposed experiment would consist of checking the possible violation of Bell’s inequality between two identical but independent systems under appropriate conditions. Beyond its theoretical interest, entanglement is a key topic in quantum computing and quantum technologies, so any attempt to gain a deeper understanding of it could be useful.
@@ I Introduction
Then a strict and undeniable violation of Bell’s inequality was given by Alain Aspect’s experiments aspect1 ; aspect2 where each single channel polarizer was replaced by a fast switching device redirecting the incident light beam to a differently oriented polarizer. The detection events on opposite sides of the experimental apparatus were separated by a space-like interval, thus free of locality loopholes. Further confirmations of the violation of Bell’s inequality were then obtained at increasingly large spatial separations nl1 ; nl2 ; nlz ; nl3 ; nl4 . Therefore, quantum nonlocality is undeniably a property of our physical world, but Bohr’s viewpoint can sound tautologic and unsatisfactory compared to our description of physical phenomena based on events located in space and evolving in time, so that the existence of correlations between events separated by an arbitrarily large space-like interval is a conundrum that goes against our intuitive perception of the physical world. Of course, not only quantum mechanics compels us to give up our intuitive perception of the physical world, nevertheless we can wonder whether - to some extent - we can recover a ”causal” description of quantum nonlocality by paying a different price with respect to Bohr’s claim of indivisibility of quantum systems.
The price to pay in the framework proposed in the present work is to consider the existence of an extra time dimension. In what follows there is no pretence of proposing something that could qualify as a new theory, rather, by means of a toy model we want to move some first steps in exploring whether nonlocality can be given a ”more familiar” interpretation by considering that while performing the measurement of an entangled state this appears as actually composed of independently existing units that exchange information via a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ that propagates at finite velocity through an extra dimension of temporal kind $\tau$. All this while respecting quantum contextuality, the property of quantum phenomenology proved by the Bell-Kochen-Specker theorem BKS ; KS . This idea can be tested against an experiment that, if proved feasible and produced the hypothesized result, would make worthwhile and necessary to go beyond the toy model discussed below.
@@ I.1 A quick digression on extra time-dimensions
On the other hand, any formulation of fundamental physics with multiple times is somehow non-trivial. In fact, as clearly summarized in Ref.B3 , naive attempts to add extra time dimensions to existing frameworks have led to two major kinds of discouraging and seemingly unavoidable difficulties: the appearance of ghosts and violations of causality. Ghosts are quantum states of systems occurring with negative probability. Causality violations are exemplified by the so called ”grandfather paradox”: two-dimensions of time by allowing time travels would make possible to kill one’s ancestors before having being born, thus entailing an absurdity. This notwithstanding, an extensive work (a survey of which can be found in Ref.bars4 ) by Itzhak Bars and co-workers on a new gauge (symplectic) symmetry, called $Sp(2,{R})$, overcomes the mentioned problems and uniquely leads to the formalism of 2T-physics. In this two time theory the additional time dimension, being treated as a ”gauge” and implied by the non-triviality of the new gauge principle, seems to some extent unphysical.
@@
The paper is organized as follows. In Section II.1 we sketch the Bohm-Bub theory, then in Section II.2 we propose an extension of this theory to simultaneous measurements of a Bell state. In Section II.3 we suggest that the fundamental idea proposed in the present work could be tested against an experiment. Section III contains some concluding remarks. Finally, in the Appendix it is sketchily shown how the formulation of the above mentioned exotic KKTs can be borrowed to define a warped metric of the (3,2)-dimensional spacetime yielding the confinement of massive particles in the extra time dimension, thus explaining why we do not experience an infinitely large extra time dimension. In other words, the extra time dimension is assumed to be a ”hidden coordinate”, that is, not observable even though physically existing. This prevents the appearance of ghosts and causality violations.
@@ II Nonlinear equations for wavefunction collapse
As is well known, two postulates of quantum mechanics are somewhat conflicting because on the one side the time evolution of the state vector of a given system is described by a linear, unitary operator and, on the other side, the result of the measure of an observable projects the system into the subspace relative to the eigenvalue/eigenstate of the measured observable according to the result obtained. Thus, measuring an observable entails the so-called collapse of the state vector (or wavefunction collapse), a discontinuous breach of the unitary evolution of a quantum system, actually a nonlinear time evolution. This topic has given rise to several discussions and interpretations (see for instance Refs.bassi ; RMP ; dorje ), however, for our purpose we resort to the proposal described in the following section.
@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter our main hypothesis enters the game. We assume that the wavefunctions collapses of the two particles - that now have their own individuality - are correlated via an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real hidden physical field ${\mathscr{X}}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. We tentatively assume that the dynamical evolution of the scalar field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the field ${\mathscr{X}}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the individual outcomes of measurements on entangled entities unpredictable, outcomes which however remain perfectly correlated. In other words, although it is demonstrated that the Bell inequality is violated by non-product states [43], nevertheless, also with the assumption done in the present work that a Bell state factorizes when both entangled entities interact simultaneously with the respective measurement devices, the physical information about the entanglement - and thus the violation of Bell inequality - is maintained by the collapse equations entailing correlated outcomes of the measures via the field ${\mathscr{X}}({\bf x},t,\tau)$.
@@ II.3 Proposal for a nonconventional experiment
The source term ${\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]$ in equation (20), even if of unspecified analytic form, amounts to assuming that the field ${\mathscr{X}}({\bf x},t,\tau)$ encodes the information about the wavefunction collapse due to the measurement performed on a given system, and equation (20) tells that this information is spread in every spatial direction. Therefore, a collapse information could also reach another identical system driving its collapse and mimicking entanglement correlation. This is the hypothesis to be experimentally proved or disproved that also suggests how to design an appropriate experiment. In fact, let us now consider two identical EPR sources, schematically depicted in Figure 3, simultaneously emitting each a pair of entangled photons or particles. Alice and Bob perform measurements at the ends of the first system, and Eve and Tom at the ends of the second system. Of course, the photons (or particles with spin) reaching Alice and Bob or Eve and Tom are in state superposition and constrained to fulfil a conservation law (momentum, energy, angular momentum) which is not the case of two particles reaching, say, Alice and Eve because a-priori the entities (say photons, as in Figure 3) reaching Alice and Eve are independent, that is, not entangled. Now, as soon as the photon detected by Alice has ”chosen” its polarization state how does the photon being detected by Bob ”know” that it has no choice left for its polarization state? The orthodox reply is that they are part of the same indivisible system even if they are very far apart, but if we assume that the information about the outcome of the measurement performed by Alice is broadcasted in every spatial direction (through the extra temporal dimension), a particle belonging to an identical system, like the one reaching Eve, could be driven to collapse into the same complementary state as in Bob’s case. Therefore, let us imagine an experimental setup where the measurement apparatuses of Alice and Eve are spatially close to one another and both of them much farther from the apparatuses of their companions Bob and Tom, so that the extra-time interval $\Delta_{1}\tau$ needed by the hypothetical field ${\mathscr{X}}({\bf x},t,\tau)$ to exchange information between Alice and Bob on one side, and Eve and Tom on the other side, is much longer than the interval $\Delta_{2}\tau$ needed to exchange information between Alice and Eve. Now, the global state vector of the two independent subsystems - Alice, Bob and Eve, Tom - is the product $|\Psi_{A,B}\rangle|\Psi_{E,T}\rangle$ where
$|\Psi_{A,B}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|\phi^{(A)}_{+}\rangle|\phi^{(B)}_{-}\rangle-|\phi^{(A)}_{-}\rangle|\phi^{(B)}_{+}\rangle)$ (29)
$|\Psi_{E,T}\rangle$ $=$ $\frac{1}{\sqrt{2}}(|\phi^{(E)}_{+}\rangle|\phi^{(T)}_{-}\rangle-|\phi^{(E)}_{-}\rangle|\phi^{(T)}_{+}\rangle)\ .$ (30)
If the four measuring devices in the setup of Fig.3 simultaneously detect the photons, or particles, then the state vector $|\Psi_{A,B}\rangle|\Psi_{E,T}\rangle$ factorizes to $|\Phi^{(A)}\rangle|\Phi^{(B)}\rangle|\Phi^{(E)}\rangle|\Phi^{(T)}\rangle$ [with the notations of (12))] out of which - after the above given argument and experimental setup - correlated behaviours are expected within the state $|\Phi^{(A)}\rangle|\Phi^{(E)}\rangle$. Consequently, the central hypothesis of the present work boils down to rewriting the collapse equations (II.2) as
@@
A clarification needs to be made here. Equations (II.2) and (II.3) describe the evolution of wavefunctions collapse in ordinary time after $t_{0}$, the instant of particles interaction with the measuring devices, and ”after” a mutual exchange of information in the extra time. In the case of the proposed new experiment, the surmised phenomenon of information exchange through the extra time depends entirely on the spatial arrangement of the four measuring devices. The exchange of information between Eve and Alice (assumed at the shortest distance among the four measuring devices) could conceivably destroy the entanglement with their respective partners allowing a rearrangement of the correlations among the results of the four measuring devices. A more refined analysis would be needed to describe what kind of correlations - among all these devices - could be expected with generic spatial arrangements. Such analysis is at present beyond the scope of our toy model.
An important remark is in order. What is described in this Section aims to propose the conceptual scheme of a possible experimental verification of the physical hypothesis underlying this work on quantum entanglement. Indeed, the practical implementation of this proposal could make use of recent developments to design a more sophisticated experimental setup than the one sketched in Fig.3, resorting for example to the use of solid-state devices where the phenomenon of parametric down-conversion occurs inside a non-linear crystal generating entangled photons physrepGenovese ; mandel ; kwiat1 , as well as various methods to ensure loophole-free experiments physrepGenovese ; kwiat2 . For instance, a method for simultaneous generation of two independent pairs of entangled photons could be borrowed from Ref.zeilinger .
@@ III Concluding remarks
Let us sketchily show how we can borrow from Ref.visser - almost verbatim - the suggestion for a five dimensional space-time metric of (+ + + - -) signature - with a non-compact and infinite extra time dimension - where the particles are trapped on the standard four-dimensional space-time. Thus, by simply modifying the signature of the extra dimension of the ”exotic” Kaluza-Klein metric of Ref.visser , consider the infinitesimal arc-length in 3+2 space-time as
@@
which has the form a ultrahyperbolic equation weinstein for which a solution can be found in the form of travelling waves in the four standard directions but confined in the extra time direction $\tau$, that is
@@ Acknowledgments
The author wishes to thank Roger Penrose for an interesting and useful discussion held at the Arts Centre De Brakke Grond, Amsterdam, in 2014, a discussion that has been the remote origin of the present work. Useful comments and suggestions emerged during several discussions with Giulio Pettini, Gabriele Vezzosi, Matteo Gori, Roberto Franzosi, Guglielmo Iacomelli, and Jack Tuszynski. The author thanks Stefano Ruffo for having brought to his attention the paper in Ref.baracca . This work was partially supported by the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 964203 (FET-Open LINkS project).